# LADM: Long-context Training Data Selection with Attention-based Dependency Measurement for LLMs

### v1

@@ Abstract
Long-context modeling has drawn more and more attention in the area of Large Language Models (LLMs). Continual training with long-context data becomes the de-facto method to equip LLMs with the ability to process long inputs. However, it still remains an open challenge to measure the quality of long-context training data. To address this issue, we propose a Long-context data selection framework with Attention-based Dependency Measurement (LADM), which can efficiently identify high-quality long-context data from a large-scale, multi-domain pre-training corpus. LADM leverages the retrieval capabilities of the attention mechanism to capture contextual dependencies, ensuring a comprehensive quality measurement of long-context data. Experimental results show that our LADM framework significantly boosts the performance of LLMs on multiple long-context tasks with only 1B tokens for continual training.
@@
Jianghao Chen1,2,3,  Junhong Wu1,2, Yangyifan Xu1,2,  Jiajun Zhang1,2,4,5thanks:   Corresponding Author 1Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Zhongguancun Academy, Beijing, China 4Wuhan AI Research, 5Shanghai Artificial Intelligence Laboratory, Shanghai, China {chenjianghao2022, wujunhong2021, xuyangyifan2021}@ia.ac.cn jjzhang@nlpr.ia.ac.cn
@@ 1 Introduction
Long-context modeling for Large Language Models (LLMs) has recently drawn more and more attention. The maximum context window of current LLMs has been significantly extended to 128K tokens for GPT-4 (Achiam et al., 2023) and 1M tokens for Gemini 1.5 (Reid et al., 2024). The long-context modeling ability greatly contributes to more sophisticated applications of LLMs in various fields such as long-context retrieval (Kamradt, 2023; Xu et al., 2023; An et al., 2024b), question answering (Kočiskỳ et al., 2018; Dasigi et al., 2021; Bai et al., 2023) and summarization (Huang et al., 2021; Zhong et al., 2021; Fabbri et al., 2019).
Preparing long-context dataset and performing continual training become the de-facto framework to enrich LLMs with the ability of processing long inputs. However, quality measurement of the long-context training data has not received enough attention. Previous studies show that if the training data is composed of concatenated short samples or lacking dependencies over long contexts (Ding et al., 2024; Fu et al., 2024; Chen et al., 2024a), models may fail to learn how to handle long-range and diverse contextual dependencies. These low-quality training samples can aggravate the tendency of LLMs to ignore valuable distant contextual information, further limiting their performance on long-context tasks. Therefore, measuring the quality of long-context training data and mining high-quality training samples become crucial for enhancing the long-context modeling capability of LLMs.
@@
To address the above issue, we propose a Long-context data selection framework with Attention-based Dependency Measurement (LADM), which measures the long-context dependency with span-level attention scores. Inspired by the inherent retrieval operations of attention mechanism (Mittal et al., 2022; Wu et al., 2024), LADM aims to leverage the attention distribution to measure the relationship within long contexts. Specifically, we first train a tiny model with long-context modeling capability named Long Attention Calculator. Then, to measure the dependency of a single long-context sample, we feed the sample into the Long Attention Calculator and compute the Pairwise Focus Score (PFS) between different spans by the accumulated attention scores. Subsequently, the Aggregated Focus Score (AFS) for each span is derived by incorporating all PFS between this span and its preceding ones. Finally, we define the sample-level Contextual Dependency Score (CDS) by a weighted sum of all AFS and select samples with high CDS for continual pre-training. Experimental results show that LADM outperforms other data selection methods on various long-context tasks, achieving an average performance improvement of 2.16% for four models across different sizes and architectures on the LongBench dataset.
@@ 2.2 Long-context Training Data for LLMs
Training data quality is crucial for the long-context modeling capability of LLMs. Fu et al. (2024) focuses on domain balance and length up-sampling for long-context training data. Staniszewski et al. (2023) and Gao et al. (2024a) propose similarity-based approaches of grouping documents to construct long-context training data. He et al. (2024) designs synthesized multi-doc QA tasks and improves the long-context information searching and reflection ability. To alleviate the "lost-in-the-middle" (Liu et al., 2024) phenomenon, An et al. (2024b) introduces Information-Intensive training on long-context QA tasks and Xiong et al. (2024) designs key-value retrieval tasks for finetuning.
Our approach characterizes the contextual dependencies within long-context samples based on the attention mechanism and effectively identifies high-quality long-context data from a large pre-training corpus. Concurrently to our work, Chen et al. (2024a) proposes a framework ProLong that divides a long-context sample into segments and calculates the delta perplexity scores between individual segments as Dependency Strength without the original context. However, we capture dependencies within the complete context, providing a more comprehensive view of the contextual relationships. Moreover, our experiments are conducted on a more diverse pre-training corpus, proving the robustness and applicability of our method.
@@ 3.1 Problem Formulation
The long-context modeling capability of LLMs is obtained by continual pre-training on long-context samples. Given a dataset $\mathcal{D}$ of long-context samples with different quality levels, our objective is to select a high-quality subset $\mathcal{D}_{s}$. Specifically, we define a scoring function $s$ to measure the quality of each data sample $x\in\mathcal{D}$. Then we rank all the samples according to their scores $s(x)$ and select the top $N$ samples to construct the subset $\mathcal{D}_{s}$ as follows:
@@ 3.2 Preliminary Analysis
Building on this analysis, we further explore how to measure the contextual dependencies within long-context data. Wu et al. (2024) reveal that LLMs incorporate retrieval ability within the attention mechanism which can be reflected in the attention distribution, with higher weights assigned to previous tokens more related to the current token. Inspired by this, we propose LADM framework to quantify the contextual dependencies by analyzing the attention distribution over long contexts.
@@ 3.3 LADM Framework
For each span $s_{j}$, we calculate PFS at varied intervals, including $\mathrm{PFS}(0,j)$, $\mathrm{PFS}(1,j)$, …, $\mathrm{PFS}(j-1,j)$. This aids in understanding how the influence of previous spans varies with increasing intervals, revealing both short-range and long-range dependencies. Notably, recent studies (Xiao et al., 2024; Hsieh et al., 2024) have shown that initial and recent tokens often receive disproportionate attention weights, suggesting a predominance of the beginning and local dependencies. To better emphasize long-range dependencies, we exclude scores for the first $m$ spans ($s_{0},...,s_{m-1}$) and the local $n$ spans ($s_{j-n},...,s_{j-1}$). For computational efficiency, we then select previous spans at a stride of $d$ and compute $L=\lfloor\frac{j-m-n-1}{d}\rfloor$ scores: $\mathrm{PFS}(m,j),\mathrm{PFS}(m+d,j),...,\mathrm{PFS}(m+Ld,j)$.
We apply weights to these scores based on the lengths of the intervals, thus encouraging longer-distance dependencies. Additionally, we take the variance of these scores into consideration. A higher variance indicates that $s_{j}$ exhibits more diverse dependencies within the context, reflecting a complex dependency pattern and rich structural information of a long-context sample. Therefore, we define the Aggregated Focus Score (AFS) of span $s_{j}$ as follows:
$\sigma_{j}=\sigma\left(\mathrm{PFS}(m,j),...,\mathrm{PFS}(m+Ld,j)\right)$
$\mathrm{AFS}(j)=\sigma_{j}\displaystyle\sum_{i=0}^{L}\mathrm{PFS}(m+id,j)(j-m-id)$
@@
$\mathrm{CDS}(S)=\displaystyle\sum_{j=n_{0}}^{N-1,\,j=j+d}\frac{j}{N}\cdot\mathrm{AFS}(j)$
where $n_{0}$ is the index of the first span and $d$ is the stride. We exclude the initial spans’ $\mathrm{AFS}(j<n_{0})$, as these early spans have fewer previous spans to depend on, resulting in a less informative measurement of dependencies. For computational efficiency, we calculate the score every $d$ spans.
@@ 4.2 Baselines
We compare LADM with following methods:
@@ 4.3 Implementation Details
We rank all samples with their CDS and select top-ranking samples from each data domain, maintaining the original domain distribution. All methods use 1B tokens selected from the dataset for continual training.
@@ 4.4 Evaluation Tasks and datasets
Figure 3: The "Needle-in-the-Haystack" task performance of different data selection methods. The x-axis denotes the evaluation context length and the y-axis denotes insertion depths of the "needle".
@@
We evaluate the language modeling capability of LLMs by measuring the perplexity (PPL) on real-world long-context data. We randomly select 100 samples exceeding 32K from the test split of the Proof-Pile (Azerbayev et al., 2022) dataset and calculate the average PPL across different context window sizes.
@@
Evaluation on perplexity or synthetic retrieval tasks can not truly reflect the performances of LLMs under real-world scenarios (Hu et al., 2024; Fang et al., 2024). Therefore, we select different types of tasks from LongBench (Bai et al., 2023) for further evaluation.
@@ 5 Experimental Results
We evaluate LADM with OpenLlama-3B-v2, Llama2-7B/13B and Mistral-7B-v0.1 on language modeling tasks, synthetic long-context tasks and real-world long-context tasks. We use "OL-3B", "L-7B", "L-13B" and "M-7B" to denote these models for short.
@@ 5.1 Perplexity Evaluation
Model Method Context Window Size
2K 4K 8K 16K 32K
OL-3B Random 4.663 4.027 3.395 2.882 2.583
ProLong 4.653 4.029 3.406 2.896 2.596
LADM 4.628 4.009 3.388 2.879 2.578
L-7B Random 4.318 3.719 3.135 2.660 2.375
ProLong 4.313 3.722 3.146 2.673 2.387
LADM 4.300 3.706 3.128 2.657 2.372
M-7B Random 4.284 3.668 3.086 2.626 2.349
ProLong 4.061 3.501 2.962 2.528 2.262
LADM 4.026 3.471 2.937 2.509 2.246
L-13B Random 3.986 3.473 2.956 2.531 2.270
ProLong 3.992 3.479 2.964 2.538 2.276
LADM 3.972 3.461 2.948 2.525 2.264
Table 2: Perplexity of 100 samples from the test split of the Proof-Pile dataset. Each sample is truncated to the corresponding context window size.
@@ 5.3 Real-World Tasks Evaluation
We report the experimental results on real-word long-context tasks in Table 4. For randomly sampling method, we conduct three experiments with different random sampling seeds and calculate the average performances. Our LADM framework outperforms other methods across nearly all types of tasks and all models. Specifically, our method achieves an average performance improvement of 2.16% on four models over ProLong. Notably, for Mistral-7B, LADM achieves a substantial improvement of 10.09% on single-document QA and and 4.66% on multi-document QA. These results highlight the strong capability of our models in handling real-world tasks and validate the effectiveness of our data selection framework, indicating the importance of incorporating long-range contextual dependencies into training data. Experimental results in Appendix F also show that our method enhances LLMs’ long-context capabilities while preserving short-context performance.
@@ 6.1 Training Efficiency of LADM
Model Method Tokens SD-QA MD-QA SUM CODE AVG
OL-3B Random 1B 24.72 19.87 17.39 54.95 29.23
2B 25.16 20.59 19.11 55.14 30.00
LADM 1B 26.63 20.62 18.32 55.56 30.28
L-7B Random 1B 30.13 25.80 24.83 57.92 34.67
2B 31.91 29.12 25.18 59.86 36.52
LADM 1B 32.24 31.10 25.85 62.12 37.83
M-7B Random 1B 24.08 24.63 22.51 60.85 33.02
2B 29.89 26.60 23.61 58.84 34.73
LADM 1B 33.85 29.09 24.47 60.58 37.00
L-13B Random 1B 28.23 32.02 25.67 64.75 37.67
2B 27.93 34.87 25.00 66.13 38.48
LADM 1B 29.03 35.53 25.82 65.28 38.92
Table 3: Performance comparison of random sampling and our LADM method on the LongBench dataset. "Tokens" denotes the total training tokens.
@@
We compare the performance of our method with the random sampling baseline under different training budgets. As shown in Table 3, our method surpasses the baseline in nearly all types of tasks, even with half of the training tokens. Specifically, LADM achieves slight improvements over OpenLlama-3B and Llama2-13B, exceeding the baseline by 1.31% for Llama2-7B and 2.27% for Mistral-7B. Even when the training tokens amount to 3B and 4B for random sampling, our method can still demonstrates superior performance, as illustrated in Appendix C. These results highlight that our proposed data selection framework can effectively extract high-quality data with strong contextual dependencies from large-scale pre-training corpora, thus enhancing the long-context modeling capability of LLMs while reducing training costs.
@@ 6.2 Data Selection Efficiency of LADM
We analyze the efficiency of data selection by adjusting the stride when calculating AFS and CDS. We use $d_{AFS}$ and $d_{CDS}$ to distinguish them in the following description. The parameter $d_{AFS}$ determines the number of previous spans considered when calculating AFS for each span, while $d_{CDS}$ determines the number of AFS involved in deriving CDS for an entire sample. Smaller values of stride improve the completeness of dependency measurement but increase computational complexity.
@@
We conducted experiments on a randomly selected set of 5,000 samples from our pre-training dataset under different hyper-parameter settings. We compare both the computational overhead and the impact on the sample-level CDS. We use the Pearson Correlation Coefficient to measure the consistency of CDS across different configurations. As shown in Table 5, all the coefficient values are greater than 0.7, indicating the consistent results of data selection. For computation efficiency, since we can get all PFS of the current span and its previous spans through multiplication of the query and key matrices, smaller $d_{AFS}$ does not significantly increase computation overhead. However, smaller $d_{CDS}$ requires calculating additional AFS for new spans, resulting in a notable increase in time overhead. Based on these analyses, we select the setting $d_{AFS}=d_{CDS}=4$ to minimize the computational cost. We also present the data selection efficiency of ProLong in Table 5. With comparable efficiency, LADM can achieve better performance, as shown in Section 5, demonstrating the effectiveness and applicability of our method.

### v2

@@ Abstract
Long-context modeling has drawn more and more attention in the area of Large Language Models (LLMs). Continual training with long-context data becomes the de-facto method to equip LLMs with the ability to process long inputs. However, it still remains an open challenge to measure the quality of long-context training data. To address this issue, we propose a Long-context data selection framework with Attention-based Dependency Measurement (LADM), which can efficiently identify high-quality long-context data from a large-scale, multi-domain pre-training corpus. LADM leverages the retrieval capabilities of the attention mechanism to capture contextual dependencies, ensuring a comprehensive quality measurement of long-context data. Experimental results show that our LADM framework significantly boosts the performance of LLMs on multiple long-context tasks with only 1B tokens for continual training.  [^1] Our code is available at https://github.com/ZNLP/LADM
@@
Jianghao Chen1,2,3,  Junhong Wu1,2, Yangyifan Xu1,2,  Jiajun Zhang1,2,4thanks:   Corresponding Author 1Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Zhongguancun Academy, Beijing, China 4Wuhan AI Research {chenjianghao2022, wujunhong2021, xuyangyifan2021}@ia.ac.cn jjzhang@nlpr.ia.ac.cn
@@ 1 Introduction
Long-context modeling for Large Language Models (LLMs) has recently drawn more and more attention. The maximum context window of current LLMs has been significantly extended to 128K tokens for GPT-4 (Achiam et al., 2023) and 1M tokens for Gemini 1.5 (Reid et al., 2024). The long-context modeling ability greatly contributes to more sophisticated applications of LLMs in various fields such as long-context retrieval (Kamradt, 2023; Xu et al., 2023; An et al., 2024b), question answering (Kočiskỳ et al., 2018; Dasigi et al., 2021; Bai et al., 2023), and summarization (Huang et al., 2021; Zhong et al., 2021; Fabbri et al., 2019).
Preparing long-context dataset and performing continual training has become the de-facto framework to enrich LLMs with the ability to process long inputs. However, quality measurement of the long-context training data has not received enough attention. Previous studies show that if the training data is composed of concatenated short samples or lacking dependencies over long contexts (Ding et al., 2024; Fu et al., 2024; Chen et al., 2024a), models may fail to learn how to handle long-range and diverse contextual dependencies. These low-quality training samples can aggravate the tendency of LLMs to ignore valuable distant contextual information, further limiting their performance on long-context tasks. Therefore, measuring the quality of long-context training data and mining high-quality training samples become crucial for enhancing the long-context modeling capability of LLMs.
@@
To address the above issue, we propose a Long-context data selection framework with Attention-based Dependency Measurement (LADM), which measures the long-context dependency with span-level attention scores. Inspired by the inherent retrieval operations of the attention mechanism (Mittal et al., 2022; Wu et al., 2024), LADM aims to leverage the attention distribution to measure the relationship within long contexts. Specifically, we first train a tiny model with long-context modeling capability named Long Attention Calculator. Then, to measure the dependency of a single long-context sample, we feed the sample into the Long Attention Calculator and compute the Pairwise Focus Score (PFS) between different spans by the accumulated attention scores. Subsequently, the Aggregated Focus Score (AFS) for each span is derived by incorporating all PFS between this span and its preceding ones. Finally, we define the sample-level Contextual Dependency Score (CDS) by a weighted sum of all AFS and select samples with high CDS for continual pre-training. Experimental results show that LADM outperforms other data selection methods on various long-context tasks, achieving an average performance improvement of 2.16% for four models across different sizes and architectures on the LongBench dataset.
@@ 2.2 Long-context Training Data for LLMs
Training data quality is crucial for the long-context modeling capability of LLMs. Fu et al. (2024) focuses on domain balance and length up-sampling for long-context training data. Staniszewski et al. (2023) and Gao et al. (2024a) propose similarity-based approaches of grouping documents to construct long-context training data. He et al. (2024) designs synthesized multi-doc QA tasks and improves the long-context information searching and reflection ability. To alleviate the "lost-in-the-middle" (Liu et al., 2024) phenomenon, An et al. (2024b) introduces Information-Intensive training on long-context QA tasks and Xiong et al. (2024) designs key-value retrieval tasks for fine-tuning.
Our approach characterizes the contextual dependencies within long-context samples based on the attention mechanism and effectively identifies high-quality long-context data from a large pre-training corpus. Concurrently with our work, Chen et al. (2024a) proposes a framework ProLong that divides a long-context sample into segments and calculates the delta perplexity scores between individual segments as Dependency Strength without the original context. However, we capture dependencies within the complete context, providing a more comprehensive view of the contextual relationships. Moreover, our experiments are conducted on a more diverse pre-training corpus, proving the robustness and applicability of our method.
@@ 3.1 Problem Formulation
The long-context modeling capability of LLMs is obtained by continual pre-training on long-context samples. Given a dataset $\mathcal{D}$ of long-context samples with different quality levels, we aim to select a high-quality subset $\mathcal{D}_{s}$. Specifically, we define a scoring function $s$ to measure the quality of each data sample $x\in\mathcal{D}$. Then we rank all the samples according to their scores $s(x)$ and select the top $N$ samples to construct the subset $\mathcal{D}_{s}$ as follows:
@@ 3.2 Preliminary Analysis
Building on this analysis, we further explore how to measure the contextual dependencies within long-context data. Wu et al. (2024) reveal that LLMs incorporate retrieval ability within the attention mechanism, which can be reflected in the attention distribution, with higher weights assigned to previous tokens more related to the current token. Inspired by this, we propose LADM framework to quantify the contextual dependencies by analyzing the attention distribution over long contexts.
@@ 3.3 LADM Framework
For each span $s_{j}$, we calculate PFS at varied intervals, including $\mathrm{PFS}(0,j)$, $\mathrm{PFS}(1,j)$, …, $\mathrm{PFS}(j-1,j)$. This aids in understanding how the influence of previous spans varies with increasing intervals, revealing both short-range and long-range dependencies. Notably, recent studies (Xiao et al., 2024; Hsieh et al., 2024) have shown that initial and recent tokens often receive disproportionate attention weights, suggesting a predominance of the beginning and local dependencies. Therefore, we exclude scores for the first $m$ spans ($s_{0},...,s_{m-1}$) and the local $n$ spans ($s_{j-n},...,s_{j-1}$). We select previous spans at a stride of $d$ for further computational efficiency.
We apply weights to these scores based on the lengths of the intervals, thus encouraging longer-distance dependencies. Additionally, we incorporate the variance of these scores. A higher variance indicates that $s_{j}$ exhibits more diverse dependencies from its previous context, reflecting a complex dependency pattern and rich structural information within the long-context sample. Therefore, we define the Aggregated Focus Score (AFS) of span $s_{j}$ as follows (omitting stride sampling for clarity):
$\sigma_{j}=\sigma\left(\mathrm{PFS}(m,j),...,\mathrm{PFS}(j-n-1,j)\right)$
$\mathrm{AFS}(j)=\sigma_{j}\displaystyle\sum_{i=m}^{j-n-1}\frac{j-i}{N}\cdot\mathrm{PFS}(i,j)$
@@
$\mathrm{CDS}(S)=\displaystyle\sum_{j=n_{0}}^{N-1}\frac{j}{N}\cdot\mathrm{AFS}(j)$
where $n_{0}$ is the index of the first span and $d$ is the stride. We exclude the initial spans’ $\mathrm{AFS}(j<n_{0})$, as these early spans have fewer previous spans to depend on, resulting in a less informative measurement of dependencies. For computational efficiency, we calculate AFS at a stride of $d$ (omitted from Eq.5 for clarity).
@@ 4.2 Baselines
We compare LADM with the following methods:
@@ 4.3 Implementation Details
We rank all samples with their CDS and select the top-ranking samples from each data domain, maintaining the original domain distribution. All methods use 1B tokens selected from the dataset for continual training.
@@ 4.4 Evaluation Tasks and datasets
Figure 3: The "Needle-in-the-Haystack" task performance of different data selection methods. The x-axis denotes the evaluation context length, and the y-axis denotes insertion depths of the "needle".
@@
We evaluate the language modeling capability of LLMs by measuring the perplexity (PPL) on real-world long-context data. We collect samples exceeding 32K from the test split of the Proof-Pile (Azerbayev et al., 2022) dataset and calculate the average PPL across different context window sizes.
@@
Evaluation on perplexity or synthetic retrieval tasks can not truly reflect the performance of LLMs under real-world scenarios (Hu et al., 2024; Fang et al., 2024). Therefore, we select different types of tasks from LongBench (Bai et al., 2023) for further evaluation.
@@ 5 Experimental Results
We evaluate LADM with OpenLlama-3B-v2, Llama2-7B/13B and Mistral-7B-v0.1 on language modeling tasks, synthetic long-context tasks and real-world long-context tasks. We use "OL-3B", "L-7B", "L-13B", and "M-7B" to denote these models for short.
@@ 5.3 Real-World Tasks Evaluation
We report the experimental results on real-world long-context tasks in Table 3. For the randomly sampling method, we conduct three experiments with different random sampling seeds and calculate the average performance. Our LADM framework outperforms other methods across nearly all types of tasks and all models. Specifically, our method achieves an average performance improvement of 2.16% on four models over ProLong. Notably, for Mistral-7B, LADM achieves a substantial improvement of 10.09% on single-document QA and 4.66% on multi-document QA. These results highlight the strong capability of our models in handling real-world tasks and validate the effectiveness of our data selection framework, indicating the importance of incorporating long-range contextual dependencies into training data. Experimental results in Appendix F also show that our method enhances LLMs’ long-context capabilities while preserving short-context performance.
Model Method Context Window Size
2K 4K 8K 16K 32K
OL-3B Random 4.951 4.268 3.562 3.026 2.685
ProLong 4.941 4.271 3.573 3.040 2.701
LADM 4.910 4.247 3.553 3.022 2.683
L-7B Random 4.515 3.900 3.264 2.780 2.458
ProLong 4.516 3.906 3.275 2.792 2.470
LADM 4.481 3.878 3.252 2.773 2.453
M-7B Random 4.620 3.936 3.267 2.775 2.455
ProLong 4.293 3.696 3.095 2.644 2.346
LADM 4.266 3.673 3.076 2.629 2.332
L-13B Random 4.200 3.657 3.084 2.645 2.349
ProLong 4.200 3.659 3.090 2.651 2.356
LADM 4.176 3.637 3.070 2.636 2.342
Table 2: Perplexity of long-context samples from the test split of the Proof-Pile dataset. Each sample is truncated to the corresponding context window size.
@@ 6.1 Training Efficiency of LADM
Model Method Tokens SD-QA MD-QA SUM CODE AVG
OL-3B Random 1B 24.72 19.87 17.39 54.95 29.23
2B 25.16 20.59 19.11 55.14 30.00
ProLong 1B 25.21 20.37 15.36 53.67 28.65
2B 27.38 18.32 19.20 55.26 30.04
LADM 1B 26.63 20.62 18.32 55.56 30.28
L-7B Random 1B 30.13 25.80 24.83 57.92 34.67
2B 31.91 29.12 25.18 59.86 36.52
ProLong 1B 29.50 29.04 24.46 59.04 35.51
2B 30.10 30.58 25.84 59.52 36.51
LADM 1B 32.24 31.10 25.85 62.12 37.83
M-7B Random 1B 24.08 24.63 22.51 60.85 33.02
2B 29.89 26.60 23.61 58.84 34.73
ProLong 1B 23.76 24.43 22.73 61.82 33.18
2B 27.41 28.85 24.29 63.36 35.98
LADM 1B 33.85 29.09 24.47 60.58 37.00
L-13B Random 1B 28.23 32.02 25.67 64.75 37.67
2B 27.93 34.87 25.00 66.13 38.48
ProLong 1B 28.94 33.20 25.62 64.62 38.04
2B 33.55 34.37 24.84 62.37 38.78
LADM 1B 29.03 35.53 25.82 65.28 38.92
Table 4: Performance comparison of random sampling and our LADM method on the LongBench dataset.
We compare the performance of our method with baseline methods under different training budgets. As shown in Table 4, our method surpasses the baseline even with half of the training tokens. Specifically, LADM achieves slight improvements over OpenLlama-3B and Llama2-13B, exceeding the baseline by 1.31% for Llama2-7B and 2.27% for Mistral-7B. Even when the training tokens amount to 3B and 4B for random sampling, our method can still demonstrate superior performance, as illustrated in Appendix C. These results highlight that our proposed data selection framework can effectively extract high-quality data with strong contextual dependencies from large-scale pre-training corpora, thus enhancing the long-context modeling capability of LLMs while reducing training costs.
@@ 6.2 Data Selection Efficiency of LADM
We first analyze the computational complexity of each stage in the LADM framework. Each PFS calculation has a complexity of $O(l^{2}\cdot d_{k})$ The complexity for AFS is $O\left(\frac{N\cdot l^{2}\cdot d_{k}}{d_{AFS}}\right)$ and for the overall CDS is $O\left(\frac{N^{2}\cdot l^{2}\cdot d_{k}}{d_{CDS}\cdot d_{AFS}}\right)=O\left(\frac{L^{2}\cdot d_{k}}{d_{CDF}\cdot d_{AFS}}\right)$ , where $l=128$ , $N=256$ , $L=32\text{k}$ , and $d_{AFS},d_{CDS}$ are the strides for calculating AFS and CDS. The two parameters decide the number of PFS calculations required and reduce the cost of full attention calculation with complexity $O(L^{2}\cdot d_{k})$ , thus affecting the data selection efficiency.
@@
We conducted experiments on a randomly selected set of 5,000 samples from our pre-training dataset under different hyper-parameter settings. We compare both the computational overhead and the impact on the sample-level CDS. We use the Pearson Correlation Coefficient to measure the consistency of CDS across different configurations. As shown in Table 5, all the coefficient values are greater than 0.7, indicating the consistent results of data selection. For computation efficiency, since we can get all PFS of the current span and its previous spans through multiplication of the query and key matrices, a smaller $d_{AFS}$ does not significantly increase computation overhead. However, smaller $d_{CDS}$ requires calculating additional AFS for new spans, resulting in a notable increase in time overhead. Based on these analyses, we select the setting $d_{AFS}=d_{CDS}=4$ to minimize the computational cost. We also present the data selection efficiency of ProLong in Table 5. With comparable efficiency, LADM can achieve better performance, as shown in Section 5, demonstrating the effectiveness and applicability of our method.